### v2

# Tradeoffs Between Alignment and Helpfulness in Language Models

Abstract
Language model alignment has become an important component of AI safety, allowing safe interactions between humans and language models, by enhancing desired behaviors and inhibiting undesired ones. It is often done by tuning the model or inserting preset aligning prompts. Recently, representation engineering, a method which alters the model’s behavior via changing its representations post-training, was shown to be effective in aligning LLMs (Zou et al., 2023a). Representation engineering yields gains in alignment oriented tasks such as resistance to adversarial attacks and reduction of social biases, but was also shown to cause a decrease in the ability of the model to perform basic tasks. In this paper we study the tradeoff between the increase in alignment and decrease in helpfulness of the model. We propose a theoretical framework which provides bounds for these two quantities, and demonstrate their relevance empirically. Interestingly, we find that while the helpfulness generally decreases, it does so quadratically with the norm of the representation engineering vector, while the alignment increases linearly with it, indicating a regime in which it is efficient to use representation engineering. We validate our findings empirically, and chart the boundaries to the usefulness of representation engineering for alignment.
1 Introduction
Advancements in large language model (LLM) development over the last few years have given LLMs a variety of abilities that allow them to serve as general purpose assistants in a wide range of tasks, such as broad-scoped question answering, writing assistance, teaching, and more  (Radford et al., 2019; Devlin et al., 2019; Brown et al., 2020; Schulman et al., 2023; OpenAI, 2023; Bubeck et al., 2023; Nori et al., 2023; West, 2023; Park et al., 2023a). The vast use of LLMs for such purposes has raised concerns due to the harm they can cause their users, such as feeding fake information (Lin et al., 2022; Weidinger et al., 2022), behaving offensively and feeding social biases (Hutchinson et al., 2020; Venkit et al., 2022; Weidinger et al., 2022), or encouraging problematic behaviors by users Roose (2023); Atillah (2023). Alignment is often the term given for the process of removing these undesired behaviors (Yudkowsky, 2001; Taylor et al., 2016; Amodei et al., 2016; Shalev-Shwartz et al., 2020; Hendrycks et al., 2021; Pan et al., 2022; Ngo, 2022).
There are several different approaches to performing alignment in LLMs, such as including aligning prompts Askell et al. (2021); Rae et al. (2021) which was shown to improve alignment and decrease toxicity in LLMs, and the procedure of reinforcement learning from human feedback (RLHF) which trains language models to be helpful and harmless (Bai et al., 2022). Though effective to an extent, these approaches are still dangerously frail, as several works have shown that adversarial prompts can trigger negative behaviors in LLMs  Wallace et al. (2019); Yu & Sagae (2021); Xu et al. (2021); Subhash (2023); Zou et al. (2023b) and the work of Wolf et al. (2023) provides a theoretical framework which shows that frozen LLMs can be misaligned with sufficiently long prompts.
Recently, a new method for alignment has been proposed by Zou et al. (2023a), which controls the model at the internal representation level by adding tailored vectors to the hidden layers’ representations. This is done by extracting directions in the model’s latent space that connect contrasting behaviors, in order to steer from undesired to desired behaviors. With it, they demonstrated experimentally that the procedure can significantly improve alignment, e.g., in resistance to adversarial attacks, though at the cost of somewhat reducing the helpfulness of the model. Similar methods have also been used by Jorgensen et al. (2023); Leong et al. (2023); Liu et al. (2023b); Turner et al. (2023) to improve alignment and reduce toxicity.
There are also known limitations to editing representations. The work of Elazar et al. (2021) empirically demonstrates how projecting out supervised linear probe directions via iterative nullspace projection can reduce performance on selected tasks. Understanding the tradeoff between model helpfulness and alignment is important for designing safe LLM systems that incorporate representation engineering. In this work we aim to shed light on the benefits and limitations of representation engineering for LLM alignment, i.e., how much does alignment improve and what is the cost in terms of the model’s abilities. We approach this question theoretically at first, and then provide empirical evidence for the validity of our theory.
Figure 1: Effect of representation engineering on helpfulness and alignment. Our main results show that alignment can improve at the cost of helpfulness. Moreover, we show that for small representation engineering norms the helpfulness decreases quadratically while the alignment increase is linear, so there is a regime in which representation engineering can be cost-effective.
In sections 3 and 4, we set up our theoretical framework and theoretical results respectively. We find that representation engineering increases alignment linearly with the representation engineering vector norm (theorem 1), while the helpfulness of the model, defined as the probability of answering a query correctly, decreases quadratically with the representation engineering vector norm (theorem 2). This means that when injecting vectors of small norms, the improvement of alignment is initially faster than the decrease in helpfulness, possibly indicating that there is a regime where prompt engineering is more effective, allowing for inference time alignment while maintaining the model’s helpfull capabilities. See figure 1 for an illustration of this intuition.
In section 5 we explore the validity of our assumptions and results in an experimental setting: We calculate alignment as a function of representation engineered vector norms corresponding to the desired behaviors and find that it increases as predicted by 1. Then, we calculate the average helpfulness over datasets of questions from different domains as a function of injected vector norms for the same behaviors and find that the decay with increased vector norm described in theorem 2 is manifested. Together, these results correspond to the intuitive illustration in figure 1 .
2 Related Work on Reperesentation Engineering
The so-called method of representation engineering was proposed by Zou et al. (2023a) as an alternative method for LLM alignment. Representation editing for LLM alignment purposes has been used in several other notable works: Jorgensen et al. (2023) show that using this method with the mean of the representation differences as steering vectors can more efficiently reduce toxicity, and Leong et al. (2023) also use this method for reducing toxicity. Liu et al. (2023b) combine this method with human feedback to improve model’s behaviors to human preferences. A similar method had been proposed by Turner et al. (2023) which subtracts layer activations in order to steer a model in latent space towards more desirable behaviors. These works propose steering directions by subtracting hidden states of prompts that display opposite behaviors. Thus, as they show experimentally, when added to the representation of a new query, its representation is steered in latent space towards a direction that leads to a more desirable behavior from the model output.
The method of representation editing is more veteran than the above alignment related efforts; it was previously used for style transfer (Shen et al., 2017), knowledge editing (Hernandez et al., 2023), controlled text-generation (Dathathri et al., 2019), truthfulness enhancement (Li et al., ), replacement of in-context learning prompts with in-context learning vectors (Liu et al., 2023a), and more. Theoretical works on the subject show that in context learning is equivalent to inserting a query with a task vector Hendel et al. (2023) and the vectors that connect representations of token pairs are parallel when the semantic difference between the pairs is similar (the linear representation hypothesis) Park et al. (2023b).
3 Necessary Preliminaries
We denote $P_{\theta}(\cdot|s)$ as the model’s next token probability distribution with the parameters $\theta$ to the prompt $s$. The model is composed of $L$ layers, $r_{\theta}^{l}$ is the $l$’th hidden state representation of the model. The next token prediction of a model is parametrized as:
$P_{\theta}(t_{n+1}|t_{1}...t_{n})=softmax(Ur^{(L)}_{\theta}(t_{1}...t_{n}))_{t_{n+1}}$ (1)
Where $r^{(L)}_{\theta}(s)$ is the final hidden layer’s representation of the prompt $s$ and $U$ is a matrix from the hidden state to a vocabulary of tokes. This is an accurate parametrization for state-of-the-art LLMs.
Parameterizing representation engineering is done by modifying each layer’s hidden state via adding a corresponding engineered vector: Denote a representation engineered model $\mathbb{P}_{\theta}$ of $L$ layers, with a set of engineered representations $R_{e}=(r_{e}^{(l=1)},...,r_{e}^{(l=L-1)})$ by $\mathbb{P}_{\theta,r_{e}}$, which is applied by adding to each hidden state the corresponding engineered vector:
$r_{\theta}^{(l)}\leftarrow r_{\theta}^{(l)}+r_{e}^{(l)}$ (2)
Note that here $l<L$, as used in Zou et al. (2023a). Additionally, we follow existing methods for representation engineering and provide a uniform norm for all the injected vectors $|r_{e}^{(l)}|=|r_{e}|$ (where the coefficient $r_{e}$ can be positive or negative) or $|r_{e}^{(l)}|=0$ for layers that are not injected.
We use the behavior expectation definition of alignment as in Wolf et al. (2023). Here we will use a binary behavior scoring function:
Definition 1.
Let $B:\Sigma\rightarrow\{-1,+1\}$ be a binary behavior scoring function, the behavior of a prompted model $\mathbb{P}(\cdot|q)$ is defined as:
$B[\mathbb{P}(\cdot|q)]=\mathbb{E}_{a\sim\mathbb{P}(\cdot|q)}[B(a)]=$ (3)
$=\sum_{a_{+}\in aligned}\mathbb{P}_{\theta}(a_{+}|q)-\sum_{a_{-}\in misaligned}\mathbb{P}_{\theta}(a_{-}|q)$ (4)
Notice that the behavior expectation is in the range $[-1,+1]$. In 1 we will prove that representation engineering is an effective method by lower bounding the behavior expectation. Notice that high probability of outputting a positive/negative response gives a positive/negative contribution to the behavior expectation, thus the sign and absolute value of behavior expectation is a good measure for the alignment of a model.
We define the model’s helpfulness as its ability to produce correct answers for queries. This can be measured as the likelihood of outputing a correct answer to a query. In order to test helpfulness for a specific skill, we can look at the probabiltiy for a specific query. See figure 1 for an example. The helpfulness of a model with respect to a specific query is defined as:
$helpfulness(model,q)=P_{\theta}(a_{correct}|q)$ (5)
Where $P_{\theta}(a_{correct}|q)$ is the model’s probability of outputing the correct answer $a$ to the query $q$. By this definition, the helpfulness is in the range $[0,1]$.
4 Main Results
We will show that representation engineering improves alignment, but harms helpfulness. Theorem 1 shows that behavior expectation is bounded from below by a hyperbolic tangent, such that it approaches $+1$ for increasing size of injected vectors and increases linearly within a bounded range. Theorem 2 shows that the helpfulness is maximized in the vicinity of norm zero injected vectors (i.e., no representation engineering) and in corollary 2 that as the norm is increased, the helpfulness decays to random guessing. The assumptions used to prove the theorems are presented formally in appendix A.
The following statement quantifies how alignment is improved by representation engineering.
Theorem 1.
Let $\mathbb{P}_{\theta,r_{e}}(\cdot|q)$ be a model prompted with query $q$ and injected with representations of coefficient $r_{e}$. Let $B:\Sigma^{*}\rightarrow\{-1,+1\}$ be a behavior scoring function. If the change to the final hidden layer due to $r_{e}$ classifies the representations of aligned and misaligned answers with respect to $B$, with margin $\Delta$, the behavior expectation of the model conditioned on the query $q$ satisfies:
$B[\mathbb{P}_{\theta,r_{e}}(\cdot|q)]\geq tanh(\Delta\lambda\cdot r_{e}+arctanh(B_{0}))$ (6)
Where $B_{0}=B[\mathbb{P}_{\theta}(\cdot|q)]$ is the behavior expectation without representation engineering and $\lambda$ is a model dependent coefficient relating between $r_{e}$ and the corresponding final hidden state norm.
As can be seen in the mathematical expression and in figure 2 for $B_{0}=-0.5$, this lower bound is a shifted hyperbolic tangent function w.r.t $r_{e}$. At $r_{e}=0$ the bound gives $B_{0}$, which is the unaltered model’s behavior. As $r_{e}$ is increased, the bound approaches $+1$, meaning the behavior asymptotically approaches $+1$. We see that for $B_{0}$ that is not too close $-1$, the increase in behavior expectation is linear due to the hyperbolic tangent’s nature, while if it is very close to $-1$, $r_{e}$ is to be increased before seeing the linear effect. Thus for behaviors on which the model is negative but also has a small tendency for positive answers, the linear effect should be felt near $r_{e}=0$. For proof see appendix section B.
Figure 2: Plot of the upper bound on behavior expectation in theorem 1
Note that while the theorem takes into account the change to the final hidden layer representation, this change incorporates the changes due to the steering vectors from all previous layers, $R_{e}=\{r_{e}^{(1)},...,r_{e}^{(L-1)}\}$.
Furthermore, the classification condition can be softened to an imperfect classifier, as discussed in appendix A and shown in appendix in E, to yield similar results.
Now, we shall bound from above the helpfulness of the model as a function of representation engineering.
Theorem 2.
Let $\mathbb{P}_{\theta,r_{e}}(\cdot|q)$ be a model prompted with query $q$ and injected with representations of coefficient $r_{e}$. The helpfulness of the model on the query is bounded with probability $1-\frac{2}{|V|}$ by:
$\mathbb{P}_{\theta,r_{e}}(a_{correct}|q)\leq\frac{P_{0}}{P_{0}+(1-P_{0})\cdot\alpha(1+\frac{\beta^{2}}{2}r_{e}^{2})}$ (7)
Where $P_{0}=\mathbb{P}_{\theta,r_{e}=0}(\cdot|q)$ is the probability of answering correctly without representation engineering, $|V|$ is the vocabulary size and $\alpha,\beta>0$ that depend on the query.
The above bound is illustrated in figure 3 for different values of $\beta$. As can be seen, around $r_{e}=0$, the bound is parabolic, i.e.  the decrease is proportional to $-r_{e}^{2}$, this can be obtained by expanding the bound near $r_{e}=0$. On the other hand, for large $r_{e}$, we see a decay to zero at a rate proportional to $r_{e}^{-2}$, this can be obtained by expanding the bound for large $r_{e}$.
Importantly, this demonstrates that for large $r_{e}$, the helpfulness decays to zero, hence representation engineering significantly harms the model’s overall performance, while for small $r_{e}$, it can initially decrease more slowly (parabolically) around $r_{e}=0$, hence the model’s performance is relatively unharmed. For the second statement to be feasible, the true helpfulness and the bound need to be close when no representation engineering is performed, such that the parabolic decrease of the upper bound forces the helpfulness to decrease as well around $r_{e}=0$. Generally, the difference between the two at $r_{e}=0$ is bounded by $1-P_{0}$, such that for queries with high probability of being answered correctly without representation engineering, i.e. $P_{0}\approx 1$, the true helpfulness and the bound will be close, guaranteeing the parabolic bound to be meaningful.
The parameter $\alpha\in[0,1]$ measures the tightness of the bound at $r_{e}=0$, since the true helpfulness at $r_{e}=0$ is $P_{0}$, while our helpfulness bound is $\frac{P_{0}}{P_{0}+\alpha(1-P_{0})}$. Thus $\alpha=1$ means the bound at $r_{e}=0$ coincides with the true helpfulness, while smaller $\alpha$ means the bound overshoots the true helpfulness. In our results, we obtain $\alpha\leq 0.5$. Figure 3 depicts this overshooting for $\alpha=0.25$. Even so, as explained above, the tightness is at least $1-P_{0}$ regardless of $\alpha$, so it is always meaningful for queries the model is initially helpful on.
The parameter $\beta$ is a measure for the rate/curvature of the quadratic decay, as it is the coefficient multiplying $r_{e}^{2}$. It is the minimum between two weighted sums of half normal distributions, whose variance $\sigma^{2}$ depends on the dataset. In section A.3, we empirically find that $\sigma\approx 0.5$, such that the mean of each weighted sum is $\approx 0.4$ (mean of a half normal distribution), therefor $\beta$ is not too small and the decay should be felt at coefficients $r_{e}$ of order of size $1$.
Figure 3: Plot of helpfulness bound with given parameters of $P_{0}$, $\alpha$ and $\beta$.
From this, we can derive a bound on the expectation value of helpfulness on a dataset:
Corollary 1.
Under the conditions of theorem 2, the expectation value of helpfulness on a dataset is bounded by:
$\mathbb{E}_{q\in dataset}[\mathbb{P}_{\theta,r_{e}}(a_{correct}|q)]\leq\textit{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ (8)
 $\textit{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}(1-\frac{2}{|V|})\frac{1}{1+\alpha(1+\frac{\beta^{2}}{2}r_{e}^{2})}+\frac{2}{|V|}$ (9)
For $\alpha,\beta>0$ that depend on the dataset.
So for datasets that have high helpfulness expectation value (about $0.9$ or higher), this bound shows helpfulness decay quadratically.
Lastly, when considering the average helpfulness over a dataset in a scenario where the number of answers is constant, $N$ (such as multiple choice questions), we obtain that on average, the model will converge to answering $1/N$ of the questions correctly as representation engineering is increased:
Corollary 2.
Under assumption 3, the expected value of the helpfulness on a dataset of queries, $\mathbb{E}_{q\in dataset}[\mathbb{P}_{\theta,r_{e}}(a_{correct}|q)]$ is asymptotically bounded from above by $\frac{1}{N}$ as $|r_{e}|\rightarrow\infty$. Where $N$ is the number of possible answers for each query.
Intuitively, for large $|r_{e}|$, the model is uniformly random, so it will guess the correct answer with probability $\frac{1}{N}$. This can be seen in section 5.
4.1 Interpretation of results – a tradeoff between alignment and usefulness
The combination of the two results show that alignment improves linearly with the norm of the vectors injected in representation engineering while helpfulness is decreased quadratically. This means that when injecting vectors of small norms, the improvement of alignment is initially faster than the decrease in helpfulness, possibly indicating that there is a regime where prompt engineering is more effective. See figure 1 for an illustration of this intuition.
5 Empirical results
Here we will calculate alignment and helpfulness as defined above and observe how it changes as we increase the vector norms of representation engineering. The experiments were performed with Llama 2 13B chat and Llama 2 13B Touvron et al. (2023). For the aligned Llama 2 13B chat, we calculate the behavior expectation w.r.t.behaviors "harmfulness" and "racism", as a function of representation engineering vector coefficients using vectors for the corresponding behaviors. For the misaligned Llama 2 13B, we do the same w.r.t."harmlessness". Then, we calculate the probability of answering queries correctly when the model is injected with the same behavior altering vectors. Additional experimental details can be found in appendix D.
We follow the work of Zou et al. (2023a) to extract the vectors used in representation engineering: We use datasets comprised of pairs of positive and negative statements w.r.t.a behavior. The difference between the representations of the pairs are used to find latent space directions that can steer the model’s responses from negative to positive behaviors or vice versa. For the "harmfulness" behavior on Llama 2 13B chat, we extracted harmful and unharmful instructions from AdvBench Robey et al. (2021; 2022) and shareGPT respectively. For "racism" we used biased and unbiased statements from the StereoSet dataset Nadeem et al. (2020). The vectors obtained here were used to calculate behavior expectation and helpfulness of the model as the norm of the vectors increased.
To calculate the racism behavior expectation, we used a behavior scoring function that assigns an answer $B(answer)=\pm 1$ to agreeing/disagreeing with a racist statement, and calculated the expectation value of this function w.r.t.the model distribution, which amounts to the difference between probabilities of agreeing and disagreeing with a racist statement. To calculate harmfulness behavior expectation, we used the behavior scoring function that assigns an answer $B(answer)=\pm 1$ if the model agrees/refuses to answer a harmful instruction and calculated its expectation value, which amounts to the difference between probabilities of agreeing and disagreeing to answer.
Figure 4(a) shows racism behavior expectation as a function of bias PCA vector coefficients injected into the model. Figure 4(b) shows harmfulness behavior expectation as a function of harmful PCA vector coefficients injected into the model. Overall we see that on both behaviors, the behavior expectation changes like a hyperbolic tangent, as expected of theorem 1.
Figure 4: (a) Racism behavior expectation of Llama 2 13B chat as a function of coefficient of injected bias PCA vectors. (b) Harmful behavior expectation of Llama 2 13B as a function of coefficient of injected harmful PCA vectors.
Note that the behavior scoring functions defined here are positive for behaviors that are negative. This was done because the model used is already aligned, so we demonstrate a flip in its behavior from positive to negative, while in Zou et al. (2023a) they show a flip in the opposite direction. In principle, theorem 1 shows the dynamics of behavior flipping its signs due to representation engineering, thus to demonstrate it, it suffices to flip the behavior in one direction.
Still, we also perform the reversed experiment, of aligning a pretrained model, Llama 2 13B w.r.t."harmlessness" with representation engineering. To this end, we extracted harmful instructions from AdvBench Robey et al. (2021; 2022) and created pairs of instruction+harmful/unharmful responses. The vectors obtained here were used to calculate behavior expectation and helpfulness of the model as the norm of the vectors increased.
To calculate harmfulness behavior expectation, we used the behavior scoring function that assigns an answer $B(answer)=\pm 1$ if the model refuses/agrees to answer a harmful instruction and calculated its expectation value, which amounts to the difference between the difference between probabilities of disagreeing and agreeing to answer.
Figure 5 shows harmlessness behavior expectation as a function of bias PCA vector coefficients injected into the model. As can be seen, the behavior expectation changes as a hyperbolic tangent, as expected of theorem 1.
Figure 5: Harmlessness behavior expectation of Llama 2 13B as a function of coefficient of injected harmful PCA vectors.
To calculate helpfulness, we queried the model with multiple choice questions from the MMLU dataset Hendrycks et al. (2020) and calculated the probability that the model assigns the correct one. This was measured as a function of injected vector coefficients inserted to the model for the behaviors above. Figure 6a shows this for the "fairness" behavior vectors and 6b shows this for the "harmfulness" behavior vectors. As can be seen, the probability to answer correctly starts by first dropping slowly then decays to zero, similarly to the proposed bound in theorem 2. The average helpfulness is usually parabolic around $r_{e}=0$ but sometimes has a plateau instead, which corresponds to the helpfulness for different queries not being peaked exactly at $r_{e}=0$, for them the bound is not tight, hence a decrease that is not parabolic around $r_{e}=0$ can be observed. Still, the slow decrease of helpfulness for small coefficients of representation engineering is seen in those cases as well.
In figure 7, we perform the same experiment when restricting the probabilities to the answers of the multiple choice question, A,B,C,D. The same result of theorem 2 is observed. Moreover, we see that the average helpfulness at the tails saturates at $\approx 1/4$, according to corollary 2.
Figure 6: Probability of correct answer of Llama-2-13B-chat on various tests from MMLU, as function of the injected vector norm. Here we used the PCA reading vectors extracted from the fairness dataset (a) and harmful instruction dataset (b).
Figure 7: Probability of correct answer of Llama-2-13B-chat on various tests from MMLU, as function of the injected vector norm. Here we used the PCA reading vectors extracted from the fairness dataset (a) and harmful instruction dataset (b). The probability of correct answer was measured relative to the answers A, B, C, D.
Similarly, we calculate the helpfulness for Llama 2 13B as a function of coefficient of the PCA vectors used to align the model for harmlessness. Figure 8(a) shows that the probability to answer correctly starts by dropping slowly then decays to zero, similarly to the proposed bound in theorem 2. Figure 8(b) shows this when restricting the answers to "A", "B", "C", "D".
Figure 8: Probability of correct answer of Llama-2-13B on various tests from MMLU, as function of the injected vector norm. Here we used the PCA reading vectors extracted to align the model for harmlessness. (a) On entire vocabulary (b) Answers restricted to "A", "B", "C", "D".
6 Discussion
We show the dynamics of representation editing. On the one hand a linear increase in the desired behavior and on the other hand a parabolic decrease on general performance of basic tasks.
There are two limitations to the terms "linear alignment" and "quadratic helpfulness". Notice that since behavior expectation is bounded between $-1$ to $+1$, the change happens within a bounded range of coefficients $r_{e}$. It may include $r_{e}=0$ (figure 4a) but it may also be outside its range 4b), such that initially one might have to work harder before seeing the payoff. As explained in section 4, the linear increase will be seen already at $r_{e}=0$ if the model’s behavior is negative but has a small probability for positive behavior as well. Also note that the quadratic bound for helpfulness is an upper bound and as explained may not be tight, indeed the tails observed in figure 6 are more typical of exponential decay than parabolic decay. Still, the interesting observation of a helpfulness decay around $r_{e}=0$ which is initially slower than alignment increase is observed.
Notice that our framework uses the model’s next token prediction to determine alignment and helpfulness. Initially this may seem to restrict the generality of our results, but note that for alignment, it is often the first token that determines whether an answer is aligned or not ("Sorry, as an AI I cannot…" vs "Sure! Here’s…"). For helpfulness this point is more subtle, as some tasks are more intricate, however, as demonstrated by our experiments, there are numerous tasks that can be reduced to next token prediction.
One may argue that these results are applicable only when one behavior vertical is enhanced with representation engineering, and the rest remain untouched leading to a reduction in performance on them. Indeed, one may apply representation engineering in several behavior verticals to improve on many behaviors, however so long as an important vertical is left out, our results imply that the LLM’s performance on it will reduce. Furthermore, one can attempt to apply this framework for analyzing the effects of normal finetuneing on alignment and helpfulness, as it too could be summed up to a change in the representations of the LLM to maximize the likelihood of desired outputs. However, the changes of a finetuning process to a model’s representations are different than representation engineering’s change: While finetuning creates small changes to the model’s representation at each training step on several different behaviors, that sums up to a large overall change, representation engineering takes a large step in one behavioral direction, disregarding other behavior verticals. As a result, the change to the representations in a representation engineering process on one behavior creates random noise on the others (assumption 3), unlike a finetuning process where this does not necessarily happen. However, representation engineering does enjoy the benefit of an online controllable step size in the desired behavior which allows to effectively manipulate the specific behavior at inference time.
Lastly, a limitation of our results for alignment guarantee is that they may rely on injecting vectors of large norms to the model. However, for very large norms, the model becomes unhelpful to the point of outputting meaningless words.
Overall, we hope that our theoretical work will shed light on the mechanism of representation engineering, which provides a new interesting direction for language model alignment.
Acknowledgments
This research was supported by the ERC (European Research Council) and the ISF (Israel Science Foundation).
Appendix A Assumptions
In A.1 we introduce our assumptions used in proving theorems 1 and 2. We discuss them in A.2 and provide experiments to check their validity in A.3
A.1 Introduction of assumptions
We assume that for small coefficients of representation engineering $r_{e}$, the norm of the change to the last hidden layer representation is linear in $r_{e}$:
Assumption 1.
Let $\mathbb{P}_{\theta,r_{e}}(\cdot|q)$ be a language model prompted with query $q$. The change to the last hidden layer representation due to representation engineering with coefficient $r_{e}$, denoted by $\delta r_{e}(q)=r^{(L)}(q,r_{e})-r^{(L)}(q,0)$ satisfies:
$|\delta r_{e}(q)|=\lambda|r_{e}|$ (10)
For a constant $\lambda>0$ that is query dependent.
It is used in theorems 1 and 2, to relate the change to the last hidden layer to the coefficients of injected representations.
A representation of an answer to a query is defined as the latent space embedding of the answer’s token, $U^{T}e_{token}$, where $e_{i}$ is the one-hot vector of the token $i$ and $U$ is the matrix from the last layer’s hidden dimension to the vocabulary. We assume that the change to the last hidden layer of the model due to representation engineering SVM classifies representations of positive and negative answers to a query with margin $\Delta$:
Assumption 2.
Given a query $q$, the change to the last hidden layer of a model due to representation engineering, $\delta r_{e}(q)=r^{(L)}(q,r_{e})-r^{(L)}(q,0)$ , SVM classifies the representations of positive and negative answers to a query $q$ with margin $\Delta$, where the positive and negative answers are defined with respect to a behavior scoring function $B:\Sigma\rightarrow\{-1,+1\}$:
$min_{i:B(i)>0,j:B(j)<0}\bigg{\{}\big{\langle}\frac{\delta r_{e}(q)}{|\delta r_{e}(q)|},U^{T}e_{i}-U^{T}e_{j}\big{\rangle}\bigg{\}}>\Delta$ (11)
That is to say, that on the axis defined by $\delta r_{e}(q)$, positive and negative representations can be separated, and the minimal distance between representations of positive and negative answers on it is $\Delta$. It is used in theorm 1, to obtain that the probability of the aligned answers increases w.r.t.the misaligned answers as the coefficients of the injected representations increases.
Note that the above assumption can be relaxed from a hard margin to a soft margin assumption, where $\delta r_{e}(q)$ classifies the representations of positive and negative answers, but part of the misaligned/aligned answers’ representations are misclassified as aligned/misaligned. This yields similar results to theorem 1 that are shown in appendix E.
For queries whose topic is unrelated to the behavior with respect to which representation engineering is performed, we expect the change to the last layer representation to be somewhat random, since the injected vectors take behavior related queries and shift them to positive behavior, but queries that are unrelated to behaviors can be shifted in random directions: Intuitively, the correction to the final layer representation has no preference for a specific token over the correct answer token, so an incorrect answer is just as likely to be on one side or the other of the plane defined by the vertical $\delta r_{e}(q)$ that passes through the correct answer representation.
Assumption 3.
When sampling a query $q$ that is unrelated to the behavior of representation engineering, the vector $\delta r_{e}(q)=r^{(L)}(q,r_{e})-r^{(L)}(q,0)$, i.e., the resulting change to the last hidden layer representation due to the steering vectors from all layers, is random with the following coordinate-wise distribution:
$\langle\frac{\delta r_{e}(q)}{|\delta r_{e}(q)|},U^{T}e_{i}\rangle\sim N(0,\sigma^{2})$ (12)
It is used in theorem 2 to formalize that representation engineering is a "perpendicular" direction to the query’s relevant answer representations.
A.2 Discussion of assumptions
Linear last hidden layer change (assumption 1): Intuitively, when adding vectors of relatively small norms to each layer, the first order Taylor expansion with respect to the vectors is good, and it scales linearly with the coefficients of the vectors. We observe experimentally in subsection A.3 that for small coefficients, the change is indeed approximately linear. Note that it suffices to assume $|\delta r_{e}(q)|$ grows monotonically with $|r_{e}|$, but for simplicity and due to experimental observations we assume the linear dependence.
SVM classification with margin $\Delta$ (assumption 2): We expect the representation engineered vectors $r_{e}$ to be good classifiers because they are obtained by methods of finding directions in the latent space that maximize the distance between representations of positive and negative textual statements. For example, in Zou et al. (2023a) the first principle component is used as a steering vector, obtained via $pca_{1}=argmax_{v}\{\mathbb{E}_{good,bad}[\langle v,r_{good}-r_{bad}\rangle]\}$ and in Jorgensen et al. (2023) the steering vector is obtained as the average of difference between positive and negative statements $\frac{1}{N}\sum_{i=1}^{N}(r^{i}_{good}-r^{i}_{bad})$. In these examples, $r_{good}$ and $r_{bad}$ are representations of queries and not the latent space embedding of the answers, as in the definition of $\Delta$-representation-separability, but we expect the steering vectors to behave similarly on them. In subsection A.3, we show that indeed $\delta r_{e}(q)$ clusters positive and negative responses to harmful queries in the model’s latent space. In appendix E we also formulate a theorem equivalent to theorem 1, but with an imperfect classifier.
Normal token logit change (assumption 3): Random vectors in high dimension tend to have a normal distribution around a sphere. If we take an isotropic normal distribution on a sphere and look at the coordinate-wise marginal probability, then it too will look Gaussian. The change to logits does not need to be exactly normal for theorem 2 to hold, it suffices that on average the correct answer logit does not increase relative to the incorrect answers’ logits (i.e.  some incorrect answer logits are enhanced w.r.t.the correct answer logit due to representation engineering and some are reduced), but we choose this assumption due to the theoretical motivation and as can be seen in A.3, this assumption is approximately satisfied in real LLMs.
A.3 Experiments for assumptions
Here we empirically check the validity of our assumptions. The experiments were performed on Llama 2 13B chat.
Norm of final hidden layer change is linear in injected vectors
We show that $\delta r_{e}(q)$ increases linearly for coefficients that are not too large (figure 9). Here we use the above mentioned fairness PCA vectors. Here we average on different datasets from MMLU:
Figure 9: Linear increase in the norm of $U\delta r_{e}(q)$ for small coefficients.
Normal assumption
We checked that the change in logit distribution is normal (figure 10) as proposed in assumption 3 and verify that indeed it is approximately normal with the standard deviation growing linearly with the injected vector norm. Note that the latter condition is because in assumption 3, we normalize $\delta r_{e}(q)$ before taking the inner product in equation 12, while here we do not, thus we expect to see the standard deviation of the Gaussians grow linearly with $|\delta r_{e}|$ which as shown above grows linearly with $r_{e}$. In figure 10(a) and (b) we show the density of change in the top ten token’s logits due to representation engineering on the dataset "international law" from MMLU for the PCA vectors of fairness and harmfulness with different coefficients. In figure 10(c) and (d) we show that the standard deviation of the distributions obtained in (a) and (b) increases linearly with the size of the coefficient. Notice that here we show $\langle\delta r_{e}(q),U^{T}(e_{i}-e_{correct})\rangle$ distributes normally instead of $\langle\delta r_{e}(q),U^{T}e_{i}\rangle$, since if the latter distributes normally, then the former is a subtraction of two normal distributions and should also be normal. The benefit of using the former instead of the latter is that it shows explicitly that the correct answer logit change is not always on one side of the Gaussian tail, which can lead to an increase in the correct answer’s probability as $|\delta r_{e}(q)|$ is increased.
Figure 10: Distribution of the change in token logits minus the logit of the correct answer. As can be seen, it is approximately normal, and the standard deviation grows linearly with the coefficient size, which is linear in $|\delta r_{e}(q)|$.
Clustering of positive and negative answers to harmful queries
Here we aim to estimate how well $\Delta$-representation-separability (definition 2) works in practice. The condition is equivalent to:
$\langle\delta r_{e}(q),U^{T}(e_{good}-e_{bad})\rangle\geq|\delta r_{e}(q)|\cdot\Delta$ (13)
In figure 11 we plot the center of representation clusters for positive and negative answers to harmful queries as the norm of harmful vectors is increased. As can be seen, the distance between the clusters increases, which corresponds to an increase in $\mathbb{E}[\langle\delta r_{e}(q),U^{T}(e_{good}-e_{bad})\rangle]$. We can define a range of coefficients in which the increase is bounded from below by a linear curve, meaning that the change in the model’s representation separates the positive and negative answer representations, similarly to the definition of $\Delta$-representation separability, but with mean instead of min.
Figure 11: Separation between representation clusters of positive and negative behavior tokens for two harmful instructions from the AdvBench dataset. Note that the model is aligned at $r_{e}=0$, thus the common tokens of $r_{e}=0$ are of aligned answers.
Appendix B Proof of theorem 1
The behavior expectation is:
$B[\mathbb{P}_{\theta,r_{e}}(\cdot|q)]=\frac{\sum_{a_{+}\in good}\mathbb{P}_{\theta,r_{e}}(a_{+}|q)-\sum_{a_{-}\in bad}\mathbb{P}_{\theta,r_{e}}(a_{-}|q)}{\sum_{a_{+}\in good}\mathbb{P}_{\theta,r_{e}}(a_{+}|q)+\sum_{a_{-}\in bad}\mathbb{P}_{\theta,r_{e}}(a_{-}|q)}=$ (14)
$=\frac{\sum_{a_{+}\in good}exp(\langle r(q)+\delta r(q),U^{T}e_{a_{+}}\rangle)-\sum_{a_{-}\in bad}exp(\langle r(q)+\delta r(q),U^{T}e_{a_{-}}\rangle)}{\sum_{a_{+}\in good}exp(\langle r(q)+\delta r(q),U^{T}e_{a_{+}}\rangle)+\sum_{a_{-}\in bad}exp(\langle r(q)+\delta r(q),U^{T}e_{a_{-}}\rangle)}=$ (15)
Where $r(q)$ is the final hidden layer representation and $\delta r(q)$ is the change to the last hidden layer due to representation engineering on the previous layers. $a_{+}\in good$ and $a_{-}\in bad$ denote the aligned and misaligned answers respectively.
$=\frac{1-\frac{\sum_{a_{-}\in bad}exp(\langle r(q)+\delta r(q),U^{T}e_{a_{-}}\rangle)}{\sum_{a_{+}\in good}exp(\langle r(q)+\delta r(q),U^{T}e_{a_{+}}\rangle)}}{1+\frac{\sum_{a_{-}\in bad}exp(\langle r(q)+\delta r(q),U^{T}e_{a_{-}}\rangle)}{\sum_{a_{+}\in good}exp(\langle r(q)+\delta r(q),U^{T}e_{a_{+}}\rangle)}}=$ (16)
$=\frac{1-\frac{\sum_{a_{-}\in bad}exp(\langle r(q),U^{T}e_{a_{-}}\rangle)exp(\langle\delta r(q),U^{T}e_{a_{-}}\rangle)}{\sum_{a_{+}\in good}exp(\langle r(q),U^{T}e_{a_{+}}\rangle)exp(\langle\delta r(q),U^{T}e_{a_{+}}\rangle)}}{1+\frac{\sum_{a_{-}\in bad}exp(\langle r(q),U^{T}e_{a_{-}}\rangle)exp(\langle\delta r(q),U^{T}e_{a_{-}}\rangle)}{\sum_{a_{+}\in good}exp(\langle r(q),U^{T}e_{a_{+}}\rangle)exp(\langle\delta r(q),U^{T}e_{a_{+}}\rangle)}}=$ (17)
Let us look at the fraction that appears in the numerator and denominator:
$\frac{\sum_{a_{-}\in bad}exp(\langle r(q),U^{T}e_{a_{-}}\rangle)exp(\langle\delta r(q),U^{T}e_{a_{-}}\rangle)}{\sum_{a_{+}\in good}exp(\langle r(q),U^{T}e_{a_{+}}\rangle)exp(\langle\delta r(q),U^{T}e_{a_{+}}\rangle)}<$ (18)
$<\frac{\sum_{a_{-}\in bad}exp(\langle r(q),U^{T}e_{a_{-}}\rangle)\cdot\min_{a^{\prime}_{-}\in bad}\{exp(\langle\delta r(q),U^{T}e_{a^{\prime}_{-}}\rangle)\}}{\sum_{a_{+}\in good}exp(\langle r(q),U^{T}e_{a_{+}}\rangle)\cdot\max_{a^{\prime}_{+}\in good}exp(\langle\delta r(q),U^{T}e_{a^{\prime}_{+}}\rangle)}=$ (19)
$=\frac{\sum_{a_{-}\in bad}exp(\langle r(q),U^{T}e_{a_{-}}\rangle)}{\sum_{a_{+}\in good}exp(\langle r(q),U^{T}e_{a_{+}}-U^{T}e_{a_{-}}\rangle)}\cdot\frac{1}{\max_{a^{\prime}_{+}\in good,a_{-}\in bad}exp(\langle\delta r(q),U^{T}e_{a^{\prime}_{+}}\rangle)}$ (20)
$=\frac{\sum_{a_{-}\in bad}exp(\langle r(q),U^{T}e_{a_{-}}\rangle)}{\sum_{a_{+}\in good}exp(\langle r(q),U^{T}e_{a_{+}}-U^{T}e_{a_{-}}\rangle)}\cdot\frac{1}{exp(\max_{a^{\prime}_{+}\in good,a_{-}\in bad}\langle\frac{\delta r(q)}{|\delta r(q)|},U^{T}e_{a^{\prime}_{+}}\rangle\cdot|\delta r(q)|)}$ (21)
From $\Delta$ margin SVM classification of $\{U^{T}a_{+}\}_{a_{+}\in good}$ and $\{U^{T}a_{-}\}_{a_{-}\in good}$ by $\frac{\delta r(q)}{|\delta r(q)|}$ (assumption 2):
$<\frac{\sum_{a_{-}\in bad}exp(\langle r(q),U^{T}e_{a_{-}}\rangle)}{\sum_{a_{+}\in good}exp(\langle r(q),U^{T}e_{a_{+}}-U^{T}e_{a_{-}}\rangle)}\cdot\frac{1}{exp(\Delta|\delta r|)}$ (22)
Plugging this back in to the behavior expectation, we obtain:
$B[\mathbb{P}_{\theta,r_{e}}(\cdot|q)]>\frac{1-\frac{\sum_{a_{-}\in bad}exp(\langle r(q),U^{T}e_{a_{-}}\rangle)}{\sum_{a_{+}\in good}exp(\langle r(q),U^{T}e_{a_{+}}-U^{T}e_{a_{-}}\rangle)}\cdot\frac{1}{exp(\Delta|\delta r|)}}{1+\frac{\sum_{a_{-}\in bad}exp(\langle r(q),U^{T}e_{a_{-}}\rangle)}{\sum_{a_{+}\in good}exp(\langle r(q),U^{T}e_{a_{+}}-U^{T}e_{a_{-}}\rangle)}\cdot\frac{1}{exp(\Delta|\delta r|)}}=$ (23)
$=\frac{1-\frac{\sum_{a_{-}\in bad}\mathbb{P}_{\theta}(a_{-}|q)}{\sum_{a_{+}\in good}\mathbb{P}_{\theta}(a_{+}|q)}exp(-\Delta|\delta r|)}{1+\frac{\sum_{a_{-}\in bad}\mathbb{P}_{\theta}(a_{-}|q)}{\sum_{a_{+}\in good}\mathbb{P}_{\theta}(a_{+}|q)}exp(-\Delta|\delta r|)}$ (24)
$=tanh(\frac{\Delta|\delta r|-\ln(\frac{\sum_{a_{-}\in bad}\mathbb{P}_{\theta}(a_{-}|q)}{\sum_{a_{+}\in good}\mathbb{P}_{\theta}(a_{+}|q)})}{2})$ (25)
Then, notice that:
$\frac{\sum_{a_{-}\in bad}\mathbb{P}_{\theta}(a_{-}|q)}{\sum_{a_{+}\in good}\mathbb{P}_{\theta}(a_{+}|q)}=\frac{1-B_{0}}{1+B_{0}}$ (26)
Where $B_{0}=B[\mathbb{P}_{\theta}(\cdot|q)]$, and that:
$arctanh(B_{0})=-\frac{1}{2}\ln\frac{1-B_{0}}{1+B_{0}}$ (27)
Thus we obtain:
$B[\mathbb{P}_{\theta,r_{e}}(\cdot|q)]>tanh(\frac{\Delta|\delta r(q)|}{2}+arctanh(B_{0}))$ (28)
Lastly, note that for coefficients that are not too large, $|\delta r(q)|$ is proportional to the injected vector coefficient $r_{e}$ (assumption 1), hence:
$B[\mathbb{P}_{\theta,r_{e}}(\cdot|q)]>tanh(\frac{\Delta\lambda}{2}\cdot r_{e}+arctanh(B_{0}))$ (29)
Where $\lambda$ is the coefficient relating $r_{e}$ to $|\delta r(q)|$.
Appendix C Proof of theorem 2
Notice that:
$\mathbb{P}_{\theta,r_{e}}(a_{correct}|q)=\frac{\mathbb{P}_{\theta,r_{e}}(a_{correct}|q)}{1}=\frac{\mathbb{P}_{\theta,r_{e}}(a_{correct}|q)}{\mathbb{P}_{\theta,r_{e}}(a_{correct}|q)+\sum_{i\in incorrect}\mathbb{P}_{\theta,r_{e}}(a_{i}|q)}=$ (30)
$=\frac{\mathbb{P}_{\theta}(a_{correct}|q)}{\mathbb{P}_{\theta}(a_{correct}|q)+\sum_{i\in incorrect}\mathbb{P}_{\theta}(a_{i}|q)e^{\langle\delta r_{e}(q),U^{T}(e_{i}-e_{correct}(q))\rangle}}\leq$ (31)
Denote $X_{i}=\langle\frac{\delta r_{e}(q)}{|\delta r_{e}(q)|},U^{T}e_{i}\rangle$ and by $\mathbb{P}_{correct}^{0}$ the probability of answering correctly without representation engineering:
$=\frac{\mathbb{P}^{0}_{correct}}{\mathbb{P}^{0}_{correct}+\sum_{i\in incorrect}\mathbb{P}_{\theta}(a_{i}|q)e^{|\delta r_{e}(q)|(X_{i}-X_{correct})}}\leq$ (32)
Denote by $I_{\pm}=\{i\in incorrect|\pm X_{i}-X_{correct}>0\}$. Using the AM-GM inequality:
$\leq\frac{\mathbb{P}^{0}_{correct}}{\mathbb{P}^{0}_{correct}+\sum_{i\in I_{+}}\mathbb{P}_{i}^{0}e^{\frac{\sum_{i\in I_{+}}\mathbb{P}_{i}^{0}(X_{i}-X_{correct})}{\sum_{i\in I_{+}}\mathbb{P}_{i}^{0}}|\delta r_{e}(q)|}+\sum_{i\in I_{-}}\mathbb{P}_{i}^{0}e^{\frac{\sum_{i\in I_{-}}\mathbb{P}_{i}^{0}(X_{i}-X_{correct})}{\sum_{i\in I_{-}}\mathbb{P}_{i}^{0}}|\delta r_{e}(q)|}}$ (33)
Denote by ±$\mathbb{P}_{\pm}=\sum_{i\in I\pm}\mathbb{P}_{i}^{0}$ and $c_{\pm}=\frac{\sum_{i\in I_{\pm}}\mathbb{P}_{i}^{0}(X_{i}-X_{correct})}{\sum_{i\in I_{\pm}}\mathbb{P}_{i}^{0}}$. We get:
$=\frac{\mathbb{P}^{0}_{correct}}{\mathbb{P}^{0}_{correct}+\mathbb{P}_{+}e^{c_{+}|\delta r_{e}(q)|}+\mathbb{P}_{-}e^{c_{-}|\delta r_{e}(q)|}}$ (34)
$\leq\frac{\mathbb{P}^{0}_{correct}}{\mathbb{P}^{0}_{correct}+\min\{\mathbb{P}_{-},\mathbb{P}_{+}\}(e^{c_{+}|\delta r_{e}(q)|}+e^{c_{-}|\delta r_{e}(q)|})}$ (35)
$\leq\frac{\mathbb{P}^{0}_{correct}}{\mathbb{P}^{0}_{correct}+\min\{\mathbb{P}_{-},\mathbb{P}_{+}\}(1+\frac{1}{2}\min\{c_{-},c_{+}\}^{2}|\delta r_{e}(q)|^{2})}$ (36)
Lastly, note that for coefficients that are not too large, $|\delta r(q)|$ is proportional to the injected vector coefficient $r_{e}$ (assumption 1), hence:
$\leq\frac{\mathbb{P}^{0}_{correct}}{\mathbb{P}^{0}_{correct}+\min\{\mathbb{P}_{-},\mathbb{P}_{+}\}(1+\frac{1}{2}\min\{c_{-},c_{+}\}^{2}|r_{e}|^{2})}$ (37)
Under the assumption that $X_{i}$ distribute normally (assumption 3), $c_{\pm}$ are a weighted sum of positive/negative half-normal distributions. Notice that with probability $\frac{1}{|V|}$ the set $I_{\pm}$ is empty, therefor with probability $1-\frac{2}{|V|}$ both sets are not empty, thus ±$\mathbb{P}_{\pm}>0$ and $c_{+}>0$, $c_{-}<0$.
Notice that $P(X_{correct}>X_{i})=\frac{1}{2}$, thus $i\in I_{+}$ with probability $\frac{1}{2}$. Therefor $\mathbb{P}_{+}$ is a weighted sum of Bernoulli variables with weights $\{\mathbb{P}^{0}_{i}\}_{i\in incorrect}$.
Appendix D Experimental Details
D.1 Obtaining the engineered representations
We use the code provided by Zou et al. (2023a) to perform representation engineering on the Llama 2 13B chat and Llama 2 13B models Touvron et al. (2023). For creating bias vectors for Llama 2 13B chat, we use the StereoSet bias dataset Nadeem et al. (2020), create pairs of prompts "[INST]Consider the amount of bias in the following scenario:[STATEMENT][/INST]", where the statement is taken from the bias dataset (a pair consists of a prompt with an unbiased statement and a prompt with a biased statement). Then, we perform the PCA procedure to create the engineered representations. For the harmfulness vectors for Llama 2 13B chat, we do the same, but with different pairs of prompts, we take harmful instructions from AdvBench dataset Robey et al. (2021; 2022) and unharmful instructions from shareGPT, create pairs of prompts with the format "[INST][INSTRUCTION][/INST]", where the instruction is taken from one of the datasets (a pair consists of two prompts, one with a harmful instruction and one with an unharmful instruction). For the harmless vectors for Llama 2 13B, we take harmful instructions from AdvBench dataset, create pairs of harmless and harmful responses to them, sampled from Llama 2 13B chat (the misaligned answers were sampled by injecting the model with harmfulness vectors and the aligned answers were sampled without representation engineering).
In all cases, we used 128 pairs of positive and negative statements. For the bias vectors, we injected the model in layers $-11$ to $-30$ and for harmfulness/harmlesness in layers $-25$ to $-33$.
D.2 Alignment experiments
To perform the experiment measuring bias, we fed the representation engineered model biased statements, with the prompt "[INST]Do you agree with the following statement? [STATEMENT][/INST]" where the statement was taken from the racist statements in the bias dataset. We extract the probabilities for the tokens "Yes" and "No" and plot $\frac{P_{yes}-P_{no}}{P_{yes}+P_{no}}$ as a function of the injected vectors’ coefficient $r_{e}$. We averaged on 10 racist statements from the StereoSet dataset.
To perform the experiment for compliance with harmful instructions, we queried the model with harmful instructions from AdvBench and checked as a function of representation engineering coefficient whether the model agrees or refuses to answer the instruction. The answers were sampled at zero temperature for each coefficient, and averaged on 20 harmful instructions for the Llama 2 13B chat experiment and 13 harmful instructions for the Llama 2 13B experiment. Note that taking the temperature to zero is equivalent to taking the representation norms to infinity, thus the hyperbolic tangent becomes a step function, and the step appears where the probability of a positive and negative response are equally likely. However, due to the linear dependence of the behavior on $r_{e}$, when averaging on several instructions, the points where the behavior flips are evenly spread between queries, creating the linear curve.
D.3 Helpfulness experiments
As a function. We evaluate the performance of a model on an MMLU dataset by feeding questions from a test to the model in the form: "[Question][A)Choice A][B) Choice B][C) Choice C][D) Choice D] The correct answer is", then calculate the probabilities for answering "A", "B", "C" and "D" and take the correct answer’s probability. We averaged the probability of the correct answer over the data set. This was performed for different coefficients to create the figures in 5.
We also used an alternative prompt: "[INST][Question][A) Choice A][B) Choice B][C) Choice C][D) Choice D][/INST] The correct answer is" to fit the format of the prompts from which the engineered representations were constructed and obtained the following figures:
Figure 12: Probability of correct answer of Llama-2-13B-chat on various tests from MMLU, as function of the injected vector norm. Here we used the PCA reading vectors extracted from the fairness dataset (a) and harmful instruction dataset (b).
Figure 13: Probability of correct answer of Llama-2-13B-chat on various tests from MMLU, as function of the injected vector norm, when restricting answers to [A,B,C,D]. Here we used the PCA reading vectors extracted from the fairness dataset (a) and harmful instruction dataset (b).
Appendix E Relaxation to Soft Margin
In the proof of theorem 1, we use the assumption that the change to the last hidden layer representation due to representation engineering SVM classifies the representations of positive and negative answers to a query with margin $\Delta$ (as explained in appendix A). We can relax this assumption by assuming that some of the negative (positive) responses’ representations, are misclassified as aligned (misaligned) answers by $\delta r_{e}(q)$, in the sense that:
$i\in aligned,j\in misaligned:~{}~{}~{}\langle\delta r_{e}(q),U^{T}(e_{i}-e_{j})\rangle\leq\Delta$ (38)
That is, the margin $\Delta$ does not hold for every pair of aligned and misaligned answers.
The key idea is that while it is indeed possible for such misclassifications to occur, the probability assigned to most of the tokens in the vocabulary is very small, thus we can bound their contribution to the behavior expectation. To this end, we define a set of misclassified responses: $\{i\in misclassified\}$ and bound the probability mass that the model assigns them by:
$\sum_{i\in misclassified}\mathbb{P}_{\theta}(i|q)<\delta\cdot\sum_{i\in aligned}\mathbb{P}_{\theta}(i|q)$ (39)
Furthermore, we bound how "deep" the misclassified negative response representations can go into the cluster of positive answer representations:
$\min_{i\in aligned,j\in misclassified}\{\langle\delta r_{e}(q),U^{T}(e_{i}-e_{j})\rangle\}>-M$ (40)
With this, the SVM classification assumption can be modified as:
Assumption 4.
Given a query $q$, the change to the last hidden layer of a model due to representation engineering, $\delta r_{e}(q)=r^{(L)}(q,r_{e})-r^{(L)}(q,0)$ , SVM classifies the representations of positive and negative answers to a query $q$ with margin $\Delta$, where the positive and negative answers are defined with respect to a behavior scoring function $B:\Sigma^{\star}\rightarrow\{-1,+1\}$:
$\min_{i\in aligned,j\in misaligned}\bigg{\{}\big{\langle}\frac{\delta r_{e}(q)}{|\delta r_{e}(q)|},U^{T}e_{i}-U^{T}e_{j}\big{\rangle}\bigg{\}}>\Delta$ (41)
Up to a set of misclassified answers, whose probability is bounded by $\sum_{i\in misclassified}\mathbb{P}_{\theta}(i|q)<\delta\cdot\sum_{i\in aligned}\mathbb{P}_{\theta}(i|q)$ that satisfy:
$\min_{i\in aligned,j\in misclassified}\{\langle\delta r_{e}(q),U^{T}(e_{i}-e_{j})\rangle\}>-M$ (42)
Note that realistically, $\delta$ can be very small for a very large set of tokens, as in inference, LLMs typically assign high probability to few tokens and very low probability for most. Hence it suffices to classify just a few high probability tokens.
We can restate theorem 1 in the following way:
Theorem 3.
Let $\delta,\epsilon>0$ and let $\mathbb{P}_{\theta,r_{e}}(\cdot|q)$ be a model prompted with query $q$ and injected with representations of coefficient $r_{e}$. Let $B:\Sigma^{\star}\rightarrow\{-1,+1\}$ be a behavior scoring function. Under assumption 4, for $r_{e}<\frac{\log\frac{\epsilon}{2\delta}}{M\cdot\lambda}$ the behavior expectation of the model conditioned on the query $q$ satisfies:
$B[\mathbb{P}_{\theta,r_{e}}(\cdot|q)]\geq tanh(\Delta\lambda\cdot r_{e}+arctanh(B_{0}))-\epsilon$ (43)
Where $B_{0}=B[\mathbb{P}_{\theta}(\cdot|q)]$ is the behavior expectation without representation engineering and $\lambda$ is a model dependent coefficient relating between $r_{e}$ and the corresponding final hidden state norm.
Proof.
We follow the proof of theorem 1, up to equation 24, there, we introduce the misclassified tokens’ contributions, which we denote by $R=\frac{\sum_{a\in misclassified}exp(\langle r(q)+\delta r_{e}(q),U^{T}e_{a})}{\sum_{a_{+}\in good}exp(\langle r(q)+\delta r_{e}(q),U^{T}e_{a_{+}})}$:
$B[\mathbb{P}_{\theta,r_{e}}(\cdot|q)]>\frac{1-\frac{\sum_{a_{-}\in bad}\mathbb{P}_{\theta}(a_{-}|q)}{\sum_{a_{+}\in good}\mathbb{P}_{\theta}(a_{+}|q)}exp(-\Delta|\delta r|)-R}{1+\frac{\sum_{a_{-}\in bad}\mathbb{P}_{\theta}(a_{-}|q)}{\sum_{a_{+}\in good}\mathbb{P}_{\theta}(a_{+}|q)}exp(-\Delta|\delta r|)+R}$ (44)
Following the same idea as with equation 18, we obtain that:
$R<\frac{\sum_{a\in misclassified}exp(\langle r(q),U^{T}e_{a})}{\sum_{a_{+}\in good}exp(\langle r(q),U^{T}e_{a_{+}})}\frac{1}{exp(-|\delta r|M)}$ (45)
Plugging this in gives:
$B[\mathbb{P}_{\theta,r_{e}}(\cdot|q)]>\frac{\sum_{a_{+}\in good}\mathbb{P}_{\theta}(a_{+}|q)-\sum_{a_{-}\in bad}\mathbb{P}_{\theta}(a_{-}|q)exp(-\Delta|\delta r|)-\sum_{a\in misclassified}\mathbb{P}_{\theta}(a|q)exp(M|\delta r|)}{\sum_{a_{+}\in good}\mathbb{P}_{\theta}(a_{+}|q)+\sum_{a_{-}\in bad}\mathbb{P}_{\theta}(a_{-}|q)exp(-\Delta|\delta r|)+\sum_{a\in misclassified}\mathbb{P}_{\theta}(a|q)exp(M|\delta r|)}>$ (46)
Denote the first second and third terms respectively as $A,B,C$:
$=\frac{A-B-C}{A+B+C}=\frac{\frac{A-B}{A+B}-\frac{C}{A+B}}{1+\frac{C}{A+B}}>(\frac{A-B}{A+B}-\frac{C}{A+B})(1-\frac{C}{A+B})>\frac{A-B}{A+B}-2\frac{C}{A+B}$ (47)
Notice that from the transition in equation 25:
$\frac{A-B}{A+B}=tanh(\frac{\Delta|\delta r|-\ln(\frac{\sum_{a_{-}\in bad}\mathbb{P}_{\theta}(a_{-}|q)}{\sum_{a_{+}\in good}\mathbb{P}_{\theta}(a_{+}|q)})}{2})$ (48)
Is the bound from theorem 1, and the second term:
$\frac{C}{A+B}=\frac{\sum_{a\in misclassified}\mathbb{P}_{\theta}(a|q)exp(M|\delta r|)}{\sum_{a_{+}\in good}\mathbb{P}_{\theta}(a_{+}|q)+\sum_{a_{-}\in bad}\mathbb{P}_{\theta}(a_{-}|q)exp(-\Delta|\delta r|)}<\delta\cdot exp(M|\delta r|)$ (49)
Lastly, notice that:
$\frac{\sum_{a_{-}\in bad}\mathbb{P}_{\theta}(a_{-}|q)}{\sum_{a_{+}\in good}\mathbb{P}_{\theta}(a_{+}|q)}=\frac{1-B_{0}}{1+B_{0}}$ (50)
Where $B_{0}=B[\mathbb{P}_{\theta}(\cdot|q)]$, and that:
$arctanh(B_{0})=-\frac{1}{2}\ln\frac{1-B_{0}}{1+B_{0}}$ (51)
Thus we obtain:
$B[\mathbb{P}_{\theta,r_{e}}(\cdot|q)]>tanh(\frac{\Delta|\delta r(q)|}{2}+arctanh(B_{0}))-2\delta\cdot exp(M|\delta r|)$ (52)
Then, note that for coefficients that are not too large, $|\delta r(q)|$ is proportional to the injected vector coefficient $r_{e}$ (assumption 1), hence:
$B[\mathbb{P}_{\theta,r_{e}}(\cdot|q)]>tanh(\frac{\Delta\lambda}{2}\cdot r_{e}+arctanh(B_{0}))-2\delta\cdot exp(M\lambda\cdot r_{e})$ (53)
Where $\lambda$ is the coefficient relating $r_{e}$ to $|\delta r(q)|$. Thus for $r_{e}<\frac{\log\frac{\epsilon}{2\delta}}{M\cdot\lambda}$:
$B[\mathbb{P}_{\theta,r_{e}}(\cdot|q)]>tanh(\frac{\Delta\lambda}{2}\cdot r_{e}+arctanh(B_{0}))-\epsilon$ (54)